OpenStudy (anonymous):

Cynthia made the conjecture that the sum of any prime number and any composite number is a composite number. Which equation is a counterexample to her conjecture? A. 11 + 2 = 13 B. 13 + 2 = 15 C. 11 + 9 = 20 D. 5 + 8 = 13

OpenStudy (wolframwizard):

To identify which example counters her conjecture, look at the parts of each example and see if they contradict what she says.

OpenStudy (anonymous):

i did that

OpenStudy (ybarrap):

11 (prime) +9 (composite) =20 (composite) Nope - not a counterexample 11 (prime) +2 (prime) =13 (prime) Nope - not a counterexample Do the same for the others. You are looking for (prime) +(composite) =(prime) or (composite) +(prime) =(prime)

OpenStudy (anonymous):

what

OpenStudy (ybarrap):

The conjecture is that a composite + prime = composite; but if you can find a case where this is not true (that is, a counterexample), then the conjecture is false. That is why you need to look for (prime) +(composite) =(prime) or (composite) +(prime) =(prime) This would then violate the conjecture.